OpenStudy (anonymous):

Which of the following is the slope between the two points (-1, -1) and (8, 0)? 9 -9 one ninth -one ninth

OpenStudy (anonymous):

slope=y2-y1/x2-x1=-1-0/-1-8 -1/-9 =1/9